# Deep Fourier-embedded Network for RGB and Thermal Salient Object Detection

### v2

@@ IV-A Experimental Setup
$\alpha=0.5$ 0.866 0.025 0.901 0.016 0.871 0.025 0.886 0.028
$\alpha=1.0$ 0.872 0.024 0.905 0.015 0.881 0.024 0.891 0.027
$\alpha=1.5$ 0.863 0.025 0.903 0.015 0.876 0.025 0.893 0.026
$\alpha=2.0$ 0.859 0.026 0.892 0.016 0.869 0.026 0.878 0.031
The S-measure [79] ($Sm$) evaluates the structural similarity between the predictions and the ground-truths. The F-measure [80] (${F_{\beta}}$) is a comprehensive metric, calculated by weighted harmonic averaging on precision and recall. We also used the weighted F-measure [81] (${wF_{\beta}}$) to complement the F-measure, addressing issues such as interpolation defects and dependency defects. Additionally, we calculated the Mean absolute error [82] ($\mathcal{M}$), which represents the average discrepancy between each pixel in the prediction and the ground-truth. Finally, Precision-recall (PR) curve [22] is drawn based on precision-recall pairs, where curves closer to $\left({1,1}\right)$ indicate better model performance.
3) Implementation Details: The backbone of DFENet was initialized with the training parameters of CDFFormer-m [66], while the remaining parameters were initialized using Kaiming initialization [83]. The model was trained on an NVIDIA Tesla P100 GPU with 16GB of memory using Adam optimization [84] and a batch size of 3. Training lasted for 200 epochs, starting with an initial learning rate of $3\times 10^{-5}$, which decayed to one-tenth every 100 epochs. We set the input size to $384\times 384$, and performed a series of data augmentation operations like random cropping, horizontal flipping, and multi-angle rotation to prevent overfitting.
The hyperparameters of CFL, specifically the scaling factor $\alpha$ in Eq. (38), were determined through careful testing. We fixed $\alpha=1$ for our experiments, unless stated otherwise. The representative values of $\alpha$ used in the tests are presented in Table I.
IV-B Model Comparison
To validate the effectiveness of the proposed DFENet, we conducted a comprehensive comparison with fifteen state-of-the-art (SOTA) RGB-T SOD models, namely CSRNet [62], MIDD [56], OSRNet [63], SwinNet [19], ADF [13], TNet [61], MGAI [14], LSNet [33], IFFNet [18], HRTNet [20], CAVER [22], WaveNet [26], LAFB [58], CAFCNet [59], and VST++ [25]. To ensure a fair comparison, all saliency maps used for the model comparison were acquired from the corresponding papers’ homepages or by running open-source code.
TABLE II: Comparison of E-measure ($Em$), S-measure ($Sm$), F-measure (${F_{\beta}}$), Weighted F-measure (${wF_{\beta}}$), and Mean Absolute Error ($\mathcal{M}$) Between Our DFENet and State-of-the-Art RGB-T SOD Models on VT821, VT1000, VT5000, and VI-RGBT1500 datasets. The Best Results are Labeled Red and the Second Best Results are Labeled Blue.
Model Backbone VT821 VT1000 VT5000 VI-RGBT1500 Param(M) FLOPs(G)
@@ IV-B Model Comparison
CSRNet21 [62] ESPNetv2 0.908 0.885 0.821 0.830 0.038 0.925 0.918 0.878 0.877 0.024 0.905 0.868 0.797 0.811 0.042 - - - - - 4.6 4.2
MIDD21 [56] VGG-16 0.895 0.871 0.760 0.804 0.045 0.933 0.915 0.856 0.822 0.027 0.897 0.868 0.763 0.801 0.043 0.870 0.762 0.665 0.765 0.069 52.4 257.9
OSRNet22 [63] VGG-16 0.896 0.875 0.801 0.813 0.043 0.935 0.926 0.891 0.892 0.022 0.908 0.875 0.807 0.823 0.040 0.917 0.872 0.811 0.836 0.044 42.4 15.6
SwinNet22 [19] Swin-B 0.926 0.904 0.818 0.847 0.030 0.947 0.938 0.894 0.896 0.018 0.942 0.912 0.846 0.865 0.026 0.948 0.901 0.863 0.882 0.031 198.7 124.3
ADF23 [13] VGG-16 0.842 0.810 0.626 0.716 0.077 0.921 0.910 0.804 0.847 0.034 0.891 0.864 0.722 0.778 0.048 - - - - - 83.1 77.6
TNet23 [61] ResNet-50 0.919 0.899 0.841 0.842 0.030 0.937 0.929 0.895 0.889 0.021 0.927 0.895 0.840 0.846 0.033 0.943 0.894 0.855 0.869 0.034 87.0 39.7
MGAI23 [14] ResNet-50 0.913 0.891 0.824 0.829 0.031 0.935 0.929 0.893 0.885 0.021 0.915 0.883 0.815 0.824 0.034 0.938 0.881 0.834 0.862 0.038 86.8 94.2
LSNet23 [33] MobileNetv2 0.911 0.878 0.809 0.825 0.033 0.935 0.925 0.887 0.885 0.023 0.915 0.877 0.806 0.825 0.037 0.935 0.883 0.835 0.859 0.037 5.4 1.2
IFFNet23 [18] Segformer 0.918 0.907 0.849 0.848 0.029 0.947 0.938 0.912 0.900 0.017 0.938 0.905 0.856 0.864 0.028 - - - - - - -
HRTNet23 [20] HRFormer 0.929 0.906 0.849 0.853 0.026 0.945 0.938 0.913 0.900 0.017 0.945 0.912 0.870 0.871 0.025 0.947 0.899 0.870 0.892 0.030 58.9 17.3
CAVER23 [22] ResNet-101 0.929 0.898 0.846 0.854 0.026 0.949 0.938 0.912 0.906 0.016 0.935 0.900 0.849 0.856 0.028 0.949 0.900 0.868 0.881 0.029 93.8 63.9
WaveNet23 [26] WaveMLP 0.929 0.912 0.863 0.857 0.024 0.952 0.945 0.921 0.909 0.015 0.940 0.912 0.865 0.867 0.026 0.951 0.909 0.874 0.879 0.027 80.7 64.0
LAFB24 [58] Res2Net-50 0.915 0.884 0.817 0.842 0.034 0.945 0.932 0.905 0.905 0.018 0.931 0.893 0.841 0.857 0.030 - - - - - 453.0 139.7
CAFCNet24 [59] ResNet-50 0.927 0.891 0.845 0.856 0.028 0.951 0.935 0.915 0.913 0.017 0.939 0.900 0.858 0.869 0.027 - - - - - - -
VST++24 [25] Swin-B 0.921 0.904 0.850 0.848 0.030 0.951 0.943 0.920 0.908 0.016 0.945 0.915 0.873 0.875 0.025 - - - - - 217.7 -
DFENet (Ours) CDFFormer-m 0.938 0.916 0.873 0.872 0.024 0.956 0.945 0.924 0.905 0.015 0.947 0.916 0.876 0.881 0.024 0.955 0.912 0.882 0.891 0.027 149.6 139.5
@@
Figure 6: F-measure curves and precision-recall (PR) curves of the proposed DFENet and fifteen existing models on the VT821, VT1000, VT5000, and VI-RGBT1500 datasets.
@@
In addition, the comparison of F-measure and PR curves in Fig. 6 further validated the leading performance of our DFENet. In the $1^{st}$ row, F-measure curves of DFENet demonstrated the highest ${F_{\beta}}$ across various thresholds. In the $2^{nd}$ row, PR curves of DFENet were closer to the point with coordinates (1,1) compared to those of other models. These findings indicate DFENet’ superior balance between precision and recall.
2) Qualitative Evaluation: Fig. 7 illustrated the saliency maps generated by our proposed DFENet and the fifteen existing models. It is evident that our DFENet can more accurately detect salient objects in various complex scenes. Even in cases where one modality fails locally ($2^{nd}$-$3^{rd}$, $6^{th}$-$7^{th}$ rows) or both modalities provide poor-quality information ($2^{nd}$ and $7^{th}$-$8^{th}$ rows), DFENet can mitigate noise and extraneous contextual cues, enabling precise delineation of salient objects from the background. This is because our proposed MPA reliably captures long-range dependencies and effectively utilizes bi-modal complementary information. Moreover, in fine-grained object detection, existing models, including WaveNet, often incur losses in object depiction. In contrast, DFENet excels in capturing intricate object details, such as the legs of a dog ($5^{th}$ row), a person on a motorcycle ($7^{th}$ row), and the frame of a chair ($10^{th}$ row). This superior performance can be attributed to the collaboration of our proposed DFENet’s components, namely FEB, FRCAB, and CFL. FEB provides reliable edge details through the depth frequency decomposition of low-level features. FRCAB enhances the model’s robustness, improving its performance in various scenes. More critically, CFL effectively addresses the problem of lost visual signals and hard frequencies unattended in the frequency domain.
Figure 7: Visual comparison of our DFENet with fifteen existing models in various complex scenarios, including large objects ($1^{st}$-$3^{rd}$, $10^{th}$ rows), small objects ($5^{th}$-$7^{th}$ rows), multiple objects ($4^{th}$ and $9^{th}$ rows), complex background ($1^{st}$-$4^{th}$, $6^{th}$ and $8^{th}$ rows), low light or contrast ($3^{rd}$ and $8^{th}$ rows), small thermal difference ($2^{nd}$-$3^{rd}$, $6^{th}$-$7^{th}$, and $9^{th}$ rows), low-quality RGB images ($2^{nd}$ and $7^{th}$ rows), and fine-grained objects ($5^{th}$, $7^{th}$, and $10^{th}$ rows). GT stands for ground-truth.
TABLE III: Ablation Study of Each Component in Our DFENet. The Best Results are Labeled Red and the Second Best are Labeled Blue.
Type Settings VT821 VT1000 VT5000 VI-RGBT1500 Params(M) FLOPs(G)
@@
Backbone w/ GFNet-H-B 0.896 0.865 0.746 0.771 0.046 0.915 0.883 0.852 0.858 0.036 0.887 0.836 0.745 0.784 0.045 0.876 0.796 0.706 0.774 0.062 136.5 129.1
w/ AFNO 0.913 0.877 0.778 0.806 0.039 0.932 0.912 0.876 0.877 0.026 0.915 0.871 0.798 0.822 0.038 0.916 0.853 0.782 0.820 0.048 132.7 138.0
w/ DFFormer-m 0.926 0.909 0.858 0.857 0.026 0.958 0.945 0.923 0.903 0.015 0.947 0.917 0.875 0.881 0.024 0.955 0.914 0.884 0.893 0.026 151.3 138.0
Module w/o MPA 0.925 0.908 0.855 0.855 0.026 0.952 0.941 0.918 0.896 0.016 0.942 0.910 0.863 0.874 0.026 0.949 0.906 0.872 0.878 0.028 144.2 137.9
w/ MHCA 0.912 0.898 0.836 0.843 0.029 0.946 0.940 0.909 0.886 0.018 0.936 0.903 0.851 0.862 0.029 0.929 0.885 0.834 0.851 0.038 146.3 138.8
w/o FRCAB 0.927 0.904 0.846 0.849 0.028 0.954 0.943 0.916 0.896 0.016 0.941 0.906 0.855 0.867 0.028 0.951 0.908 0.868 0.883 0.030 127.1 77.4
w/ RCAB 0.929 0.908 0.857 0.865 0.027 0.954 0.945 0.923 0.904 0.015 0.946 0.914 0.874 0.878 0.025 0.952 0.909 0.880 0.886 0.027 148.1 139.5
w/o FEB 0.927 0.911 0.854 0.849 0.027 0.954 0.946 0.916 0.889 0.016 0.945 0.916 0.864 0.868 0.026 0.950 0.912 0.870 0.879 0.029 148.6 135.6
w/o CFL 0.932 0.909 0.863 0.865 0.024 0.955 0.943 0.922 0.911 0.015 0.946 0.913 0.873 0.879 0.024 0.952 0.907 0.878 0.888 0.028 149.6 139.5
Input size $384^{2}$rightarrow $512^{2}$ 0.930 0.909 0.862 0.867 0.026 0.951 0.942 0.919 0.912 0.015 0.945 0.913 0.873 0.884 0.025 0.954 0.910 0.881 0.893 0.026 149.6 248.0
Full model 0.938 0.916 0.873 0.872 0.024 0.956 0.945 0.924 0.905 0.015 0.947 0.916 0.876 0.881 0.024 0.955 0.912 0.882 0.891 0.027 149.6 139.5
IV-C Ablation Study
To comprehensively assess the effectiveness of each component in our proposed DFENet, we conducted ablation studies by comparing models with (w/) and without (w/o) these components, and replacing them with classical methods that operate similarly. The quantitative and qualitative evaluations of ablation studies are shown in Table III, Fig. 8, and Fig. 9.
1) Effectiveness of backbone network: To validate the effectiveness of CDFFormer-m [66] as the backbone of DFENet, we compared it with three other FFT-based backbones, namely GFNet-H-B [64], AFNO [65], and DFFormer-m [66]. Notably, AFNO has the same dimension and depth as CDFFormer-m. As shown in the $1^{st}$-$2^{nd}$ rows of Table III, the model with GFNet-H-B and AFNO yielded unsatisfactory results, with inferior performance across all four datasets compared to the model with CDFFormer-m. This trend was further highlighted in Fig. 8, where these models struggled to accurately detect salient objects. Meanwhile, the model with DFFormer-m demonstrated uneven performance on the four datasets. Although it exhibited a slight advantage on the VI-RGBT1500 dataset ($3^{rd}$ row of Table III), it substantially lagged behind the model with CDFFormer-m on the VT821 dataset.
In addition, we attempted to apply spatial domain-based backbones to DFENet, such as VGG [85], ResNet [86], and Swin Transformer [17]. However, instability (loss reported as “Nan”) occurred during training due to gradient instability caused by the incompatibility of the spatial domain operations and the two Fourier transform operations in MPA.
2) Effectiveness of MPA (Sec. III-B): First, we removed MPA and employed element-wise addition to fuse the bi-modal features, labeled “w/o MPA” in Table III and Fig. 8. The performance of the model lacking MPA declined obviously on the four datasets due to its inability to extract complementary information. In Fig. 8, when the salient region in one or both modalities showed a different appearance, the model without MPA encountered considerable interference, impeding accurate detection of complete objects.
@@ IV-C Ablation Study
Figure 9: Visualization of the feature maps pre- and post-processed by each component. L1 to L4 indicates the layers from low to high. (a) shows RGB images, thermal images, and the feature map after FEB. (b) and (c) stand for RGB and thermal feature maps generated by the encoder. (d) represents the fused feature map obtained by MPA. (e) is the feature maps after FRCAB.
To further validate the effectiveness of the Modal-coordinated Perception Filter (MPF), the core of MPA, we replaced MPF with the widely used multi-head cross-attention (MHCA) [7, 22]. Due to the considerable complexity and memory demands of MHCA, we applied it only to the last three low-resolution layers, while the largest resolution layer still employed element-wise addition. As shown in the $5^{th}$ row of Table III and Fig. 8, the model with MHCA did not perform as well as expected, and even had less gain than the model with element-wise addition. The experiments highlight the suboptimal performance of MHCA compared to MPA in high-resolution feature fusion and demonstrate the incompatibility of spatial-domain operations with frequency-domain models.
Finally, we visualized in Fig. 9 the feature changes after processing by MPA. Even in instances of localized failure within the RGB (e.g., case 1) or thermal modalities (e.g., cases 2 and 3), MPA demonstrated remarkable resilience, successfully fusing both modalities without succumbing to noise interference. The outcomes of the above experiments collectively reinforce the high efficacy of MPA.

### v3

@@ IV-B Evaluation Metrics
$\alpha=0.5$ 0.868 0.024 0.899 0.015 0.873 0.025 0.885 0.028
$\alpha=1.0$ 0.872 0.023 0.903 0.014 0.884 0.023 0.893 0.026
$\alpha=1.5$ 0.865 0.024 0.901 0.014 0.878 0.024 0.897 0.026
$\alpha=2.0$ 0.860 0.026 0.894 0.016 0.871 0.025 0.880 0.029
Additionally, we calculated the Mean absolute error [91] ($\mathcal{M}$), which represents the average discrepancy between each pixel in the prediction and the ground-truth. Finally, Precision-recall (PR) curve [24] is drawn based on precision-recall pairs, where curves closer to $\left({1,1}\right)$ indicate better model performance.
IV-C Implementation Details
The backbone of FreqSal was initialized with the training parameters of CDFFormer-M [75], while the remaining parameters were initialized using Kaiming initialization [92]. The model was trained on an NVIDIA Tesla P100 GPU with 16GB of memory using Adam optimization [93] and a batch size of 3. Training lasted for 240 epochs, starting with an initial learning rate of $2\times 10^{-5}$, which decayed to one-tenth every 80 epochs. We set the input size to $384\times 384$, and performed a series of data augmentation operations like random cropping, horizontal flipping, and multi-angle rotation to prevent overfitting.
The hyperparameters of CFL, specifically the scaling factor $\alpha$ in Eq. (38), were determined through careful testing. We fixed $\alpha=1$ for our experiments, unless stated otherwise. The representative values of $\alpha$ are presented in Table I.
TABLE II: Comparison of E-measure (EmEm), S-measure (SmSm), F-measure ($F_{\beta}$}), Weighted F-measure ($wF_{\beta}$}), and Mean Absolute Error ($\mathcal{M}$) Between Our FreqSal and State-of-the-Art RGB-T SOD Models on the VT821, VT1000, VT5000, and VI-RGBT1500 datasets. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. FPS Was Tested on an NVIDIA RTX4060 Laptop GPU.
Model Backbone VT821 VT1000 VT5000 VI-RGBT1500 Param(M) FLOPs(G) FPS
@@ IV-C Implementation Details
CNN-based Model
CSRNet21 [69] ESPNetv2 0.908 0.885 0.821 0.830 0.038 0.925 0.918 0.878 0.877 0.024 0.905 0.868 0.797 0.811 0.042 0.836 0.715 0.572 0.661 0.089 4.6 4.2 40.2
MIDD21 [60] VGG-16 0.895 0.871 0.760 0.804 0.045 0.933 0.915 0.856 0.822 0.027 0.897 0.868 0.763 0.801 0.043 0.870 0.762 0.665 0.765 0.069 52.4 257.9 86.3
OSRNet22 [70] VGG-16 0.896 0.875 0.801 0.813 0.043 0.935 0.926 0.891 0.892 0.022 0.908 0.875 0.807 0.823 0.040 0.917 0.872 0.811 0.836 0.044 42.4 15.6 95.6
ADF23 [14] VGG-16 0.842 0.810 0.626 0.716 0.077 0.921 0.910 0.804 0.847 0.034 0.891 0.864 0.722 0.778 0.048 0.888 0.831 0.664 0.777 0.074 83.1 77.6 7.3
TNet23 [64] ResNet-50 0.919 0.899 0.841 0.842 0.030 0.937 0.929 0.895 0.889 0.021 0.927 0.895 0.840 0.846 0.033 0.943 0.894 0.855 0.869 0.034 87.0 39.7 59.0
LSNet23 [40] MobileNetv2 0.911 0.878 0.809 0.825 0.033 0.935 0.925 0.887 0.885 0.023 0.915 0.877 0.806 0.825 0.037 0.935 0.883 0.835 0.859 0.037 5.4 1.2 88.4
MGAI23 [15] ResNet-50 0.913 0.891 0.824 0.829 0.031 0.935 0.929 0.893 0.885 0.021 0.915 0.883 0.815 0.824 0.034 0.938 0.881 0.834 0.862 0.038 86.8 94.2 4.7
LAFB24 [62] Res2Net-50 0.915 0.884 0.817 0.842 0.034 0.945 0.932 0.905 0.905 0.018 0.931 0.893 0.841 0.857 0.030 0.937 0.887 0.847 0.867 0.035 453.0 139.7 29.3
Transformer-based Model
SwinNet22 [21] Swin-B 0.926 0.904 0.818 0.847 0.030 0.947 0.938 0.894 0.896 0.018 0.942 0.912 0.846 0.865 0.026 0.948 0.901 0.863 0.882 0.031 198.7 124.3 25.8
IFFNet23 [20] Segformer 0.918 0.907 0.849 0.848 0.029 0.947 0.938 0.912 0.900 0.017 0.938 0.905 0.856 0.864 0.028 - - - - - - - -
HRTNet23 [22] HRFormer 0.929 0.906 0.849 0.853 0.026 0.945 0.938 0.913 0.900 0.017 0.945 0.912 0.870 0.871 0.025 0.947 0.899 0.870 0.892 0.030 58.9 17.3 11.5
CAVER23 [24] ResNet-101 0.929 0.898 0.846 0.854 0.026 0.949 0.938 0.912 0.906 0.016 0.935 0.900 0.849 0.856 0.028 0.949 0.900 0.868 0.881 0.029 93.8 63.9 16.2
CAFCNet24 [30] ResNet-50 0.927 0.891 0.845 0.856 0.028 0.951 0.935 0.915 0.913 0.017 0.939 0.900 0.858 0.869 0.027 0.947 0.894 0.866 0.881 0.030 185.7 324.2 13.8
VST++24 [27] Swin-B 0.921 0.904 0.850 0.848 0.030 0.951 0.943 0.920 0.908 0.016 0.945 0.915 0.873 0.875 0.025 - - - - - 217.7 - -
MAGNet24 [72] SMT 0.930 0.909 0.863 0.865 0.026 0.949 0.938 0.917 0.912 0.016 0.943 0.909 0.870 0.876 0.025 0.949 0.899 0.873 0.885 0.029 16.1 9.9 28.1
PATNet24 [66] P2T-B 0.935 0.914 0.872 0.870 0.024 0.952 0.942 0.920 0.910 0.015 0.947 0.918 0.879 0.883 0.023 0.948 0.908 0.878 0.885 0.029 94.8 51.1 11.8
UniTR24 [28] Swin-B 0.936 0.902 0.863 0.875 0.025 0.958 0.939 0.925 0.926 0.014 0.950 0.909 0.878 0.892 0.023 0.951 0.895 0.874 0.898 0.029 146.3 135.0 31.0
Frequency-based Model
WaveNet23 [41] WaveMLP 0.929 0.912 0.863 0.857 0.024 0.952 0.945 0.921 0.909 0.015 0.940 0.912 0.865 0.867 0.026 0.951 0.909 0.874 0.879 0.027 80.7 64.0 5.3
FreqSal (Ours) CDFFormer-M 0.939 0.917 0.873 0.872 0.023 0.959 0.948 0.928 0.903 0.014 0.951 0.920 0.879 0.884 0.023 0.956 0.916 0.885 0.893 0.026 149.6 139.5 17.1
@@
Figure 7: F-measure curves and precision-recall (PR) curves of the proposed FreqSal and eighteen existing RGB-T SOD models on the VT821, VT1000, VT5000, and VI-RGBT1500 datasets.
IV-D Model Comparison
To validate the effectiveness of the proposed FreqSal, we conducted a comprehensive comparison with eighteen state-of-the-art (SOTA) RGB-T SOD models, namely CSRNet [69], MIDD [60], OSRNet [70], SwinNet [21], ADF [14], TNet [64], MGAI [15], LSNet [40], IFFNet [20], HRTNet [22], CAVER [24], WaveNet [41], LAFB [62], CAFCNet [30], VST++ [27], MAGNet [72], PATNet [66], and UniTR [28]. To ensure a fair comparison, all saliency maps used for the model comparison were acquired from the corresponding papers’ homepages or by running open-source code.
@@ IV-D Model Comparison
In addition, the F-measure and PR curves in Fig. 7 further validate the leading performance of our FreqSal. In the $1^{st}$ row of Fig. 7, F-measure curves of FreqSal demonstrate the highest $F_{\beta}$} across various thresholds. Moreover, PR curves of FreqSal are closer to the point with coordinates (1,1) compared to those of other models. These findings indicate FreqSal’ superior balance between precision and recall.
Figure 8: Visual comparison of our FreqSal with eighteen existing RGB-T SOD models in various complex scenarios, including large objects ($1^{st}$-$3^{rd}$, $10^{th}$ rows), small objects ($5^{th}$-$8^{th}$ rows), multiple objects ($4^{th}$ and $9^{th}$ rows), complex background ($1^{st}$-$5^{th}$, $8^{th}$ and $10^{th}$ rows), low light or contrast ($3^{rd}$-$6^{th}$ rows), small thermal difference ($2^{nd}$-$3^{rd}$, $5^{th}$, $7^{th}$, and $9^{th}$ rows), low-quality RGB images ($2^{nd}$ and $7^{th}$ rows), and fine-grained objects ($7^{th}$, and $10^{th}$ rows). GT stands for ground-truth.
2) Qualitative Evaluation: Fig. 8 illustrates the saliency maps generated by our proposed FreqSal and representative RGB-T SOD models. It is evident that our FreqSal can more accurately detect salient objects in various complex scenes. Even in cases where one modality fails locally ($2^{nd}$-$4^{th}$, $7^{th}$, and $9^{th}$ rows) or both modalities provide poor-quality information ($2^{nd}$, $5^{th}$, and $7^{th}$ rows), FreqSal can mitigate noise and extraneous contextual cues, enabling precise delineation of salient objects from the background. This is because our proposed MPA reliably captures long-range dependencies and effectively utilizes bimodal complementary information. Moreover, in fine-grained object detection, existing models, including UniTR [28], often incur losses in object depiction. In contrast, FreqSal excels in capturing intricate object details, such as the legs of a person ($2^{th}$ row), a person on a motorcycle ($7^{th}$ row), and the frame of two chairs ($10^{th}$ row). This superior performance can be attributed to the collaboration of our proposed FreqSal’s components, namely FEB, FRCAB, and CFL. FEB provides reliable edge details through the depth frequency decomposition of low-level features. FRCAB enhances the model’s robustness, improving its performance in various scenes. More critically, CFL effectively addresses the problem of lost visual signals and hard frequencies unattended in the frequency domain.
TABLE III: Ablation Study of Each Component in Our FreqSal. The Best Results are Labeled Red and the Second Best are Labeled Blue. FPS Was Tested on an NVIDIA RTX4060 Laptop GPU.
Type Settings VT821 VT1000 VT5000 VI-RGBT1500 Params(M) FLOPs(G) FPS
@@
Modality w/o Thermal 0.928 0.905 0.841 0.856 0.029 0.950 0.942 0.904 0.890 0.021 0.943 0.908 0.851 0.868 0.029 0.945 0.894 0.844 0.874 0.034 86.7 101.2 24.9
w/o RGB 0.900 0.855 0.768 0.786 0.041 0.940 0.916 0.875 0.863 0.026 0.930 0.885 0.819 0.835 0.034 0.901 0.821 0.750 0.791 0.058 86.7 101.2 24.9
Backbone w/ GFNet-H-B 0.898 0.865 0.749 0.773 0.044 0.920 0.887 0.856 0.860 0.034 0.893 0.842 0.752 0.790 0.042 0.880 0.804 0.711 0.777 0.060 136.5 129.1 19.4
w/ AFNO 0.912 0.879 0.780 0.806 0.038 0.935 0.916 0.878 0.879 0.024 0.917 0.874 0.805 0.828 0.036 0.918 0.856 0.783 0.819 0.047 132.7 138.0 7.6
w/ DFFormer-M 0.928 0.912 0.861 0.858 0.026 0.957 0.946 0.925 0.900 0.014 0.951 0.922 0.879 0.883 0.023 0.956 0.915 0.887 0.896 0.026 151.3 138.0 11.7
Module w/o MPA 0.926 0.908 0.857 0.857 0.025 0.954 0.944 0.921 0.898 0.016 0.944 0.911 0.865 0.874 0.026 0.949 0.908 0.874 0.879 0.027 144.2 137.9 18.2
w/ MHCA 0.915 0.902 0.839 0.847 0.027 0.948 0.941 0.913 0.890 0.017 0.941 0.906 0.855 0.865 0.027 0.931 0.889 0.838 0.853 0.036 146.3 138.8 16.7
w/o FRCAB 0.928 0.904 0.846 0.849 0.027 0.952 0.942 0.917 0.897 0.016 0.946 0.910 0.857 0.870 0.027 0.950 0.910 0.870 0.884 0.029 127.1 77.4 17.5
w/ RCAB 0.931 0.910 0.858 0.864 0.026 0.956 0.947 0.925 0.906 0.015 0.948 0.918 0.876 0.880 0.024 0.953 0.911 0.881 0.885 0.027 148.1 139.5 17.3
w/o FEB 0.928 0.911 0.855 0.850 0.026 0.956 0.948 0.919 0.892 0.015 0.947 0.919 0.869 0.872 0.025 0.949 0.913 0.872 0.878 0.029 148.6 135.6 17.8
w/o CFL 0.933 0.911 0.864 0.865 0.023 0.956 0.945 0.925 0.909 0.015 0.947 0.916 0.874 0.881 0.024 0.954 0.908 0.879 0.887 0.028 149.6 139.5 17.1
Full model 0.939 0.917 0.873 0.872 0.023 0.959 0.948 0.928 0.903 0.014 0.951 0.920 0.879 0.884 0.023 0.956 0.916 0.885 0.893 0.026 149.6 139.5 17.1
@@
IV-E Ablation Study
To comprehensively assess the effectiveness of each component in our proposed FreqSal, we conducted ablation studies on the VT821, VT000, VT5000, and VI-RGBT1500 datasets. We compared models with (w/) and without (w/o) these components, and replaced them with classical methods that operate similarly. The quantitative and qualitative evaluations of ablation studies are shown in Table III, Fig. 9, and Fig. 10.
1) Effectiveness of auxiliary modality: To evaluate the contribution of each auxiliary modality, we conducted experiments by removing the thermal and RGB modalities separately. As shown in the $1^{st}$ and $2^{nd}$ rows of Table III, the model’s performance declined in both cases. In particular, excluding the RGB modality led to average decreases of 3.5%, 6.1%, 9.9%, and 7.8% in $Em$, SmSm, $wF_{\beta}$}, and $F_{\beta}$}, respectively. These results confirm that both modalities are crucial for achieving accurate detection.
2) Effectiveness of backbone network: To validate the effectiveness of CDFFormer-M [75] as the backbone of FreqSal, we compared it with three other FFT-based backbones, namely GFNet-H-B [73], AFNO [74], and DFFormer-M [75]. Notably, AFNO had the same dimension and depth as CDFFormer-M. As shown in the $3^{th}$-$4^{th}$ rows of Table III, the model with GFNet-H-B and AFNO yielded unsatisfactory results, with inferior performance across all four datasets compared to the model with CDFFormer-M. This trend was further highlighted in Fig. 9, where these models struggled to accurately detect salient objects. Meanwhile, the model with DFFormer-M demonstrated uneven performance on the four datasets. Although it exhibited a slight advantage on the VI-RGBT1500 dataset ($5^{th}$ row of Table III), it substantially lagged behind the model with CDFFormer-M on the VT821 dataset.
In addition, we attempted to apply spatial domain-based backbones to FreqSal, such as VGG [94], ResNet [95], and Swin Transformer [19]. However, successive Inverse Fourier Transforms of MPA in the feature space of the spatial domain backbone lead to a quadratic amplification of the gradient, resulting in instability during training (loss reported as “NaN”).
3) Effectiveness of MPA (Sec. III-B): First, we removed MPA and employed element-wise addition to fuse the bimodal features, labeled “w/o MPA” in Table III and Fig. 9. The performance of the model lacking MPA declined obviously on the four datasets due to its inability to extract complementary information. In Fig. 9, when the salient region in one or both modalities shows a different appearance, the model without MPA encounters considerable interference, impeding accurate detection of complete objects.
Figure 10: Visualization of the feature maps pre- and post-processed by each component. L1 to L4 indicate the layers from low to high. (a) shows RGB images, thermal images, and the feature map after FEB. (b) and (c) stand for RGB and thermal feature maps generated by the encoder. (d) represents the fused feature map obtained by MPA. (e) presents the feature maps after RCAB. (f) is the feature maps after FRCAB.
To further validate the effectiveness of the MPF, the core of MPA, we replaced MPF with the widely used multi-head cross-attention (MHCA) [6, 24]. Due to the considerable complexity and memory demands of MHCA, we applied it only to the last three low-resolution layers, while the highest resolution layer still employed element-wise addition. As shown in the $7^{th}$ row of Table III and Fig. 9, the model with MHCA did not perform as well as expected, and even had less gain than the model with element-wise addition. The experiments highlight the suboptimal performance of MHCA compared to MPA in high-resolution feature fusion and demonstrate the incompatibility of spatial-domain operations with frequency-domain models.
Finally, we visualized in Fig. 10 the feature changes after processing by MPA. Even in instances of localized failure within the RGB (e.g., case 1) or thermal modalities (e.g., cases 2 and 3), MPA demonstrates remarkable resilience, successfully fusing both modalities without succumbing to noise interference. The outcomes of the above experiments collectively reinforce the high efficacy of MPA.